# Nonparametric Estimation and Comparison of Distance Distributions from Censored Data

### 26 Jan 2024 v3

@@ I Introduction
Although we are, of course, not the first to employ survival analysis for transportation geography (e.g., [12, 13]) nor the first to compare survival curves (e.g., [14, 15]), we present a framework for comparing collections of censored transportation events and provide a sampling-based U test for the task.
@@ II-B Drawing Uncensored Samples
    For each event, we construct a distance matrix $D$ with elements $D_{ij}=f(b_{\text{start}_{i}},b_{\text{end}_{j}})$ and distance function $f$. The event’s observation interval $[\min(D),\max(D)]$ includes the true distance of the transportation event. Once we have all events’ observation intervals, we sort the unique distances as $d_{1}<d_{2}<\ldots<d_{k}$ and construct a set of sorted intervals $\{(d_{1},d_{2}),(d_{2},d_{3}),\ldots,(d_{k-1},d_{k})\}$.
@@
    $d_{*}=\hat{F}^{-1}(u),\quad u\sim\text{Uniform}(0,1)$ (7)
@@
    $\epsilon\sim\mathcal{N}(0,\sigma^{2}),\quad\sigma=\frac{\hat{F}_{\text{upper},\alpha}(d_{*})-\hat{F}(d_{*})}{Q(1-\frac{\alpha}{2})}$ (8)
    $d_{\text{sample}}=d_{*}+\epsilon$ (9)
@@ II-C Stochastic Dominance
Taken together, these steps may be organized into a Monte Carlo U Test (Algorithm 1): for a predetermined number of trials, we repeatedly draw uncensored distance samples by applying Steps 1-3 of Section II-B before performing a one-sided U Test [21]. We aggregate the resultant p-values using Fisher’s method.
Algorithm 1 Monte Carlo U Test for Censored Events
@@
2:     Input: $E_{A},E_{B}$: samples of censored transportation events of types $A$ and $B$, respectively. $m,n$: numbers of samples to draw from $F_{A}$ and $F_{B}$, respectively. $n_{\text{trials}}$: number of trials to run the underlying U test.
@@
5:     Obtain $SortedInts_{A}$, $SortedInts_{B}$ from $E_{A}$ and $E_{B}$
6:     Estimate $\hat{F}_{A}(d)$ using $SortedInts_{A}$
7:     Estimate $\hat{F}_{B}(d)$ using $SortedInts_{B}$
8:     $T\leftarrow 0$
9:     for $i=1$ to $n_{\text{trials}}$ do
10:         $s_{A}\leftarrow$ $m$ samples drawn from $\hat{F}_{A}(d)$
11:         $s_{B}\leftarrow$ $n$ samples drawn from $\hat{F}_{B}(d)$
12:         $S\leftarrow$ sorted list of items from $s_{A}$ and $s_{B}$
13:         $R_{1}\leftarrow\sum_{xinS_{A}}\Big{(}\text{rank of $x$ in $S$}\Big{)}$
14:         $u=R_{1}-\frac{m(n+1)}{2}$
15:         $p_{i}\leftarrow P(U\geq u|F_{A}\sim F_{B})$
16:         $T\leftarrow T+\log(p_{i})$
17:     end for
18:     $T\leftarrow-2T$
19:     $p=P(S\geq T|T\sim\chi^{2}_{2n_{\text{trials}}})$ return $T,p$
20:end procedure
@@ II-D Statistical Distance
Alternatively, we may consider Maximum Mean Discrepancy (MMD), a statistical distance between distributions expressed as the largest difference in expected values across functions of a class $\mathcal{G}$, described in detail in [24, 25]. Absent direct access to their uncensored generating distributions, we can draw uncensored distance samples $s_{A}$ (of size $m$) and $s_{B}$ (of size $n$) from $\hat{F}_{A}$ and $\hat{F}_{B}$, respectively, and apply the empirical MMD estimate provided in [24]:
@@ III-B Simulation Results
In our trials, we see that the KS Statistic between the empirical CDF of uncensored events and the estimated CDF reconstructed from their censored realizations decreases exponentially with locale size (Figure 2A). We observe a similar effect with the empirical MMD between the uncensored events and the samples drawn from our reconstruction (Figure 2B).
We also empirically validate the calibration of the Monte Carlo U Test (Algorithm 1). Using the same simulation environment, we vary the number of locales and confirm maintenance of the Type I error rate (Figure 3) and uniformity of p-values (Table I) under the null hypothesis.
Figure 3: Monte Carlo U Test calibration results: maintenance of the Type I error rate. We consider the following numbers of locales: $10$, $50$, $100$, $500$, $1000$, $2000$. We initialize our simulation environment, generate $1000$ censored events, and run our Monte Carlo U Test with $E_{A}=E_{B}$, $m=n=100$, and $n_{\text{trials}}=100$. We run $1000$ hypothesis tests of this kind for each number of locales and illustrate the frequency of p-values below each significance level.
TABLE I: Monte Carlo U Test calibration results: uniformity of p-values.1
@@
    P-values from Figure 3 are distributed into $20$ equal-width bins, each bin covering a range of $5\%$. We report the results of chi-squared tests of uniformity under the null hypothesis.
@@
    Reproduced from Socioeconomic deprivation, travel distance, location of service, and uptake of breast cancer screening in North Derbyshire, UK by Ravi Maheswaran, Tim Pearson, Hannah Jordan, and David Black, (volume 60, page 210, copyright notice year 2006), with permission from BMJ Publishing Group Ltd., as well as the corresponding author [27].
@@
Although this is a sensible analysis, there are limitations. Because the chi-squared test treats distance as a categorical variable, the ordinality of distance is lost. Additionally, the chi-squared test aims to assess whether the expected frequency of events occurring in the distance bins differs between the ”Invited” and ”Attended” groups; even if the result is deemed significant, it does not specify the nature of the relationship. The authors also interpret distance as a continuous variable by mapping postal codes to their corresponding centroid location, a technique not uncommon in this situation, but we believe this mapping makes results difficult to interpret.
Figure 4: Kernel density estimates via samples drawn from the Invited and Attended distributions. Log-scale is used on the y-axis to highlight the tail.
We apply Algorithm 1 to the data in Table II, with the alternative hypothesis that the distribution of distances from screening center for women invited for screening stochastically dominates the distribution of distances for women actually attending screening. We assume a comfortable $100$ km upper bound on distance traveled, run $100$ trials with $100$ samples drawn per trial, and find a statistically significant result ($p<0.0001$, Figure 4). Although the modes of the distributions are very close, we see that the ”Attended” group has measurably less tail density. We can estimate an effect size via logistic regression from our drawn samples, obtaining a modest odds ratio of 0.9943 ($95\%$ CI $0.9933$, $0.9954$).
@@ IV Conclusion
In this work, we discuss the analysis of censored transportation event records. We leverage ideas from survival analysis to perform this estimation and suggest a sampling-based stochastic dominance test for comparing sets of censored transportation records. Future work should refine the approach and compare against other tests for comparing survival curves, such as the Mantel–Cox test [14].

### 24 Mar 2024 v4

@@ I Introduction
© 2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@
Although we are, of course, not the first to employ survival analysis for transportation geography (e.g., [12, 13]) nor the first to compare survival curves (e.g., [14, 15]), we present a framework for comparing collections of censored transportation events and provide a sampling-based U-test for the task.
@@ II-B Drawing Uncensored Samples
    For each event, we construct a distance matrix $D$ with elements $D_{ij}=f(b_{\text{start}_{i}},b_{\text{end}_{j}})$ and distance function $f$. The event’s observation interval $[\min(D),\max(D)]$ includes the true distance of the transportation event.
@@
    $d_{*}=\hat{F}^{-1}(u),\quad u\sim\text{Uniform}(0,1)$
@@
    $\epsilon\sim\mathcal{N}(0,\sigma^{2}),\quad\sigma=\frac{\hat{F}_{\text{upper},\alpha}(d_{*})-\hat{F}(d_{*})}{Q(1-\frac{\alpha}{2})}$
    $d_{\text{sample}}=d_{*}+\epsilon$
@@ II-C Stochastic Dominance
Taken together, these steps may be organized into a Monte Carlo U-test (Algorithm 1): for a predetermined number of trials, we repeatedly draw uncensored distance samples by applying Steps 1-3 of Section II-B before performing a one-sided U-test [21]. We aggregate the resultant p-values using Fisher’s method.
Algorithm 1 Monte Carlo U-Test for Censored Events
@@
2:     Input: $E_{A},E_{B}$: samples of censored transportation events of types $A$ and $B$, respectively. $m,n$: numbers of samples to draw from $F_{A}$ and $F_{B}$, respectively. $n_{\text{trials}}$: number of trials to run the underlying U-test.
@@
5:     Estimate $\hat{F}_{A}(d)$ using $E_{A}$
6:     Estimate $\hat{F}_{B}(d)$ using $E_{B}$
7:     $t\leftarrow 0$
8:     for $i=1$ to $n_{\text{trials}}$ do
9:         $s_{A}\leftarrow$ $m$ samples drawn from $\hat{F}_{A}(d)$
10:         $s_{B}\leftarrow$ $n$ samples drawn from $\hat{F}_{B}(d)$
11:         $S\leftarrow$ sorted list of items from $s_{A}$ and $s_{B}$
12:         $R_{1}\leftarrow\sum_{x\in S_{A}}\Big{(}\text{rank of $x$ in $S$}\Big{)}$
13:         $u=R_{1}-\frac{m(m+1)}{2}$
14:         $p_{i}\leftarrow P(U\geq u|F_{A}=F_{B})$
15:         $t\leftarrow t+\log(p_{i})$
16:     end for
17:     $t\leftarrow-2t$
18:     $p=P(T\geq t|T\sim\chi^{2}_{2n_{\text{trials}}})$ return $t,p$
19:end procedure
@@ II-D Statistical Distance
Alternatively, we may consider Maximum Mean Discrepancy (MMD), a statistical distance between distributions expressed as the largest difference in expected values across functions of a class $\mathcal{G}$, described in detail in [24, 25]. Absent direct access to their uncensored generating distributions, we can draw distance samples $s_{A}$ (of size $m$) and $s_{B}$ (of size $n$) from $\hat{F}_{A}$ and $\hat{F}_{B}$, respectively, and apply the empirical MMD estimate provided in [24]:
@@ III-B Simulation Results
In our trials, we see that the KS statistic between the empirical CDF of uncensored events and the estimated CDF reconstructed from their censored realizations decreases with locale count (Figure 2A), as expected. We observe a similar effect with the empirical MMD between the uncensored events and the samples drawn from our reconstruction (Figure 2B).
We also empirically inspect the calibration of the Monte Carlo U-test (Algorithm 1). Using the same simulation environment, we vary the number of locales and confirm maintenance of the Type I error rate (Figure 3) and uniformity of p-values (Table I) under the null hypothesis.
Figure 3: Monte Carlo U-test calibration results: maintenance of the Type I error rate. We consider the following numbers of locales: $10$, $50$, $100$, $500$, $1000$, $2000$. We initialize our simulation environment, generate $1000$ censored events, and run our Monte Carlo U-test with $E_{A}=E_{B}$, $m=n=100$, and $n_{\text{trials}}=100$. We run $1000$ hypothesis tests of this kind for each number of locales and illustrate the frequency of p-values below each significance level.
TABLE I: Monte Carlo U-test calibration results: uniformity of p-values.1
@@
    P-values from Figure 3 are distributed into $20$ equal-width bins, each bin covering a range of $5\%$. We report the results of chi-squared tests of uniformity.
@@
    Reproduced from Socioeconomic deprivation, travel distance, location of service, and uptake of breast cancer screening in North Derbyshire, UK by Ravi Maheswaran, Tim Pearson, Hannah Jordan, and David Black, (volume 60, page 210, copyright notice year 2006), with permission from BMJ Publishing Group Ltd. and the corresponding author [27].
@@
Although this is a sensible analysis, there are limitations. Because it treats distance as a categorical variable, the ordinality of distance is ignored. Additionally, the chi-squared test aims to assess whether the expected frequency of events occurring in the distance bins differs between the “Invited” and “Attended” groups; even if the result is deemed significant, it does not specify the nature of the relationship. The authors also interpret distance as a continuous variable by mapping postal codes to their corresponding centroid location, a technique not uncommon in this situation, but we believe this mapping makes results difficult to interpret.
We apply Algorithm 1 to the data in Table II, with the alternative hypothesis that the distribution of distances from screening center for women invited for screening stochastically dominates the distribution of distances for women actually attending screening. We assume a comfortable $100$ km upper bound on distance traveled, run $100$ trials with $100$ samples drawn per trial, and do not find a statistically significant result ($p=0.8813$). We repeat the procedure for “Not Attended” (i.e., those invited, but did not attend) and “Attended,” also failing to find a statistically significant result ($p=0.7671$).
@@ IV Conclusion
In this work, we discuss the analysis of censored transportation event records. We leverage ideas from survival analysis to perform this estimation and suggest a sampling-based stochastic dominance test for comparing sets of censored transportation records. Future work should refine the approach and consider other tests for comparing survival curves, such as the Mantel–Cox test [14].